x+y=3 and 3x-4y=-19

Answer by checkley77(12844)

x+y=3
x=3-y now replace x with (3-y) in the other equation.
3x-4y=-19
3(3-y)-4y=-19
9-3y-4y=-19
-7y=-19-9
-7y=-28
y=-28/-7
y=4 ans.
x+4=3
x=3-4
x=-1 ans.
Proof:
3*-1-4*4=-19
-3-16=-19
-19=-19